CAN SOME ONE HELP ME SOLVE THIS WORD PROBLEM.

Suppose that N= square root x +3 models the number of cases of an infection, in millions, of a disease x years from now.
a. approximately how many cases of the infection will there be in 9 years from now?
b. in approximatetly how many years will there be in 5 million cases?

Answer by checkley71(8403)

N=SQRT(X+3)
N=SQRT(9+3)
N=SQRT12
N=3.464 MILLION CASES IN 9 YEARS.
-----------------------------------------
5=SQRT(X+3) NOW SQUARE BOTH SIDES OF THE EQUATION:
25=X+3
X=25-3
X=22 YEARS THERE WILL BE 5,000 CASES.